OpenStudy (coconutjj):

Can someone please help: f(x) = (1+(1/x))(cos(sqrt(x^3))) f'(x) = ?

OpenStudy (anonymous):

\[f(x)=(1+\frac{1}{x})\cos(\sqrt{x^3})=(1+\frac{1}{x})\cos(x^\frac{3}{2})\]

OpenStudy (coconutjj):

yes...

OpenStudy (anonymous):

Apply product rule and use chain rule for the composite cos function

OpenStudy (coconutjj):

i'll give it a shot..

OpenStudy (coconutjj):

\[\frac{ -1 }{ x^2 }\cos (x ^{\frac{ 2 }{ 3 }})+(1+\frac{ 1 }{ x })(\sin (x ^{\frac{ 2 }{ 3 }}) * \frac{ 2 }{ 3 }x ^{\frac{ -1 }{ 3 }})\]

OpenStudy (coconutjj):

is this correct..?

OpenStudy (anonymous):

the power is 3/2 not 2/3 and there should be a minus with sin

OpenStudy (coconutjj):

oh i see

OpenStudy (coconutjj):

do i just simplify from heree

OpenStudy (anonymous):

No, you can leave it as it is, apart from a little minus sign mistake, you did pretty well, you wrote the powers wrong but if you've wrote them right Im sure you'd have done it correctly

OpenStudy (coconutjj):

thank you very much ! :)